# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

### v2

@@ 2 Related work
Programmatic reinforcement learning (PRL). PRL represents RL policies using more structured and potentially more interpretable and generalizable representations, such as decision tree (Bastani et al., 2018), state machine (Inala et al., 2020; Koul et al., 2019; Lin et al., 2024), symbolic expression (Verma et al., 2018; 2019; Bhupatiraju et al., 2018; Landajuela et al., 2021), Logic programming language (Jiang and Luo, 2019), and program written in domain-specific language (DSL) (Andre and Russell, 2001; Silver et al., 2020; Sun et al., 2020; Zhu et al., 2019; Qiu and Zhu, 2022; Moraes and Lelis, 2024; Mariño et al., 2021; Trivedi et al., 2021; Liu et al., 2023; Carvalho et al., 2024). Trivedi et al. (2021); Liu et al. (2023) devised PRL tasks in the Karel domain (Pattis, 1981) and proposed learning embedding spaces of programs using variational autoencoders (Kingma and Welling, 2014) and then optimizing program embeddings using cross-entropy method or RL. Carvalho et al. (2024) achieve the state-of-the-art results in the Karel domain (Pattis, 1981) by searching in a programmatic space of AST structure. Despite the encouraging results and the generality of these methods, they are notoriously inefficient, requiring tens of millions of program executions to obtain task-solving programs. In this work, to devise an efficient PRL framework, we make the first attempt to leverage the knowledge and the reasoning ability of LLMs to generate a set of initial programs to bootstrap search algorithms; moreover, we propose Scheduled HC, a search algorithm that best utilizes the LLM-generated programs.
Large language models for code generation. With a remarkable ability to understand and generate natural languages and codes, LLMs have been widely adopted for code generation (Brown et al., 2020; Achiam et al., 2023; Nijkamp et al., 2022; Xu et al., 2022; Roziere et al., 2023) via neural program synthesis methods (Lin et al., 2018; Sun et al., 2018; Desai et al., 2016; Raza et al., 2015; Wang et al., 2017; Le et al., 2022; Parisotto et al., 2017; Balog et al., 2017). These works target general-purpose programming languages for software development, such as Python and C++, with abundant data on the internet; in contrast, we aim to solve PRL tasks that require writing programs restricted by given domain-specific languages. Prior works exploring using LLMs to synthesize DSL programs via providing LLMs with DSL grammars and few-shot examples, hindsight relabeling, and prioritized experience replay (Wang et al., 2023a; Grand et al., 2024; Butt et al., 2024). While these works focus on string and array transformation as well as abstract reasoning (Chollet, 2019), we leverage LLMs to synthesize DSL program policies to be executed in an RL environment and maximize the return.
@@ 3 Preliminary
Programmatic space vs. latent space. In the programmatic space, a program $P$ can be represented as an abstract syntax tree (AST), where each leaf node in the AST represents a program token (Carvalho et al., 2024). To obtain its neighborhood program, a node is sampled from its AST, and then replaced by a subtree randomly generated using the DSL’s production rules and sampling strategies. Another choice is to search in a learned latent space (Trivedi et al., 2021; Liu et al., 2023), constructed by training a variational autoencoder on randomly generated DSL programs. To sample a neighbor program, a small noise is added to the latent embedding of the current program candidate, and the decoder can generate a program from the noise-corrupted embedding.
Hill climbing (HC). This search algorithm climbs whenever there is a higher place, i.e., moves to a program having a higher episodic return. Given a program, HC generates its $k$ neighborhood programs from the search space. These programs are evaluated in the environment. If one of them has a higher episodic return than the initial program, it is set as the search center program, and then this process repeats. Otherwise, the algorithm halts and returns the best program evaluated.
Cross-entropy method (CEM). CEM (Rubinstein and Kroese, 2004) first generates $k$ neighborhood programs from its latent space and evaluates their episodic return on the Karel environment (Trivedi et al., 2021). Next, if the average return over the top-$E$ neighborhood programs is higher than the best-seen average return, it continues the process on the mean latent embedding of the top-$E$ neighborhood programs. When the average return no longer increases, the program with the highest episodic return is returned.
Cross-entropy beam search (CEBS). CEBS extends CEM to maintain a set of $E$ candidate programs to perform beam search. In other words, CEBS searches all the neighborhoods of the top-$E$ programs.
Note that HC can search in both programmatic and latent spaces, while CEM and CEBS are designed for continuous search spaces, i.e., latent program space only. All of these search algorithms share the same weakness – they search surrounding programs in a narrow proximity thus the distance between the initial program and the optimal program directly upper-bounds the algorithm’s sample efficiency.
@@
4 Large language model-guided search for efficient programmatic reinforcement learning
@@ 4.2 Generating DSL programs with Pythonic-DSL strategy
Table 2: The Pythonic-DSL instruction. The Karel DSL is specified via a constrained version of Python with pre-defined functions including Pythonic-DSL perceptions that allow the agent to observe, and Pythonic-DSL actions that enable the agent to interact. This Pythonic-DSL description is fed into the LLM via the system prompt so that it can generate Python code that can later be converted into Karel DSL. We will show that this approach outperforms the direct generation of DSL or Python programs in Table 3. The full prompt is presented in Section D.1.
@@ 4.3 Optimizing the program policy with Scheduled Hill Climbing
where $n$ represents the number of evaluated programs, $k(n)$ denotes a function indicating the current number of neighborhood programs based on the number of evaluated programs $n$. This function’s logarithm is a linear combination of the logarithms of two hyperparameters, $K_{start}$ and $K_{end}$, which signify the initial and terminal numbers of neighbors to search, respectively. The variable $r(n)$ governs the linear ratio, which itself is a sinusoidal function that gradually increases from 00 to $1$ throughout the evaluation of a total of $N$ programs (see Appendix E). With this scheduler, the number of neighbors $k$ is a function of the execution budget used, growing from $K_{start}$ to $K_{end}$. We apply this scheduler to Hill Climbing and call this method Scheduled Hill Climbing (Scheduled HC).
@@ 5.1 Evaluation setup
Baselines. We compare our proposed framework with the following methods: LEAPS (Trivedi et al., 2021), HPRL (Liu et al., 2023), CEBS (Carvalho et al., 2024), and Hill Climbing (Carvalho et al., 2024) (the current state-of-the-art). More details of these methods are presented in Appendix F.
Metrics. For a fair comparison of the sample efficiency, we use the number of programs evaluated to represent the program-environment interaction budget. A method is a more sample-efficient PRL algorithm if it achieves a higher average episodic return on a fixed budget. Specifically, for each task, the number of task variances is $C$. One program evaluated means a program is executed on all $C$ task variances to obtain the average return. We set the number of task variances $C=32$, the maximum number of program evaluation $N=10^{6}$, i.e., the interaction budget. Programs achieving an average return of 1.0 will be considered optimal.
@@ 5.2 LLM-guided search significantly improves the sample efficiency
Figure 4 presents the experimental result, indicating that our proposed LLM-GS surpasses all the existing methods by a large margin on almost all the tasks. We also include a plot aggregating the performance across all ten tasks to clearly demonstrate the improvement in Appendix H (Figure 15).
On the Karel set, our framework completely solves these tasks and exhibits the best sample efficiency. Also, LLM-GS performs extraordinarily well on Seeder, OneStroke, and DoorKey, which are among the hardest of the Karel-Hard set. We highlight our DoorKey result, with an efficiency improvement from not converging at even 1M (baselines) to converging within around 50K (ours). Existing PRL algorithms struggle at DoorKey due to the fact that it is a two-stage task with a sparse reward function. Therefore, most of the baselines cannot escape the local maxima and converge around $0.5$. Curiously, our method starts at an episodic return of $\approx 0.5$ from the beginning and converges quickly without being trapped by the local maxima. To analyze, we examine one output program from LLM shown in Figure 5. The LLM clearly understands that DoorKey is a two-stage task and thus generates a corresponding two-stage structured program, which is much easier for later search to find the optimal program. The full dialogue is presented in Appendix D.
@@ 5.3 Ablation studies
To justify the effectiveness of our proposed Pythonic-DSL strategy, we conduct experiments on generating either Python or DSL programs with the same 8 seeds and present the results in Table 3. The acceptance rate calculates the ratio of executable DSL programs, while the best return refers to the highest episodic return among the legal programs. Our method achieves the highest acceptance rate on 9 / 10 tasks and the highest best return on 7 / 10 tasks. The complete prompts are presented in Appendix I.
@@
Figure 6: Comparing search algorithms. We compare the proposed Scheduled HC with latent space CEM, CEBS, HC-$250$, and Scheduled HC, and programmatic space HC-$k$$\in\{32,256,2048\}$. Our method achieves comparatively good efficiency in DoorKey and performs best in CleanHouse, justifying the efficacy of our proposed search method.
@@
Figure 7: Comparing initializations. We compare initializing the search population using LLM-generated programs or randomly sampled programs. The results show that starting from LLM-generated programs significantly improves efficiency and performance, especially in DoorKey, highlighting the effectiveness of our proposed framework.
@@ 5.4 Performance on novel tasks showcases the extensibility of LLM-GS
Figure 8: Performance of PathFollow & WallAvoider. With only changing the task description, our proposed method surpasses the best-performing baseline HC, highlighting the extensibility of LLM-GS.
A key advantage of our proposed LLM-GS framework is its ability to adapt to new tasks using only text descriptions without requiring users to have any domain knowledge or familiarity with the DSL. To showcase this feature, we designed two novel Karel tasks, PathFollow and WallAvoider, as detailed in Section C.3. By simply replacing the task descriptions while keeping the Karel domain prompt, Python-to-DSL prompt, and hyperparameters the same, we compare our method with the best-performing baseline, HC. The results presented in Figure 8 show that by only changing the task descriptions, our LLM-GS framework still achieves improved sample efficiency and finds better policies.
@@ 5.5 Beyond initialization – a preliminary attempt on LLM revision
Given the encouraging results from LLM-initialized programs, one might be curious whether LLM can progressively revise and improve its programs once given feedback from the environment, e.g., reward. For completeness, we additionally conduct studies with the following LLM-based revision:
@@
Figure 9: LLM revision. The performance gain of LLM revision saturates within a few rounds.
We conduct the LLM revision experiments on DoorKey for five revision rounds, each round with $32$ programs with $5$ seeds. The results are presented in Figure 9. We can see that the improvement quickly saturates within the first two rounds, and no significant gain is observed with more revision rounds. This indicates that only using LLM to revise the program may not be sufficient to solve the task. On the other hand, our proposed Scheduled HC can converge to an optimal program – and is “free,” i.e. no expensive API calls or “API usage has reached your notification threshold” emails. The details can be found in Appendix I.
@@ 6 Discussion
We propose an LLM-guided framework (LLM-GS), that leverages the programming skills and common sense of LLM to bootstrap the efficiency of assumption-free, random-guessing search methods. We propose a Pythonic-DSL framework to address the inability of LLMs to generate precise and grammatically correct programs. To further maximize the programs generated by LLMs, we design a search algorithm, Scheduled Hill Climbing, which can explore the programmatic search space to consistently improve the programs. The experimental results on the Karel tasks show the improved sample efficiency of our proposed LLM-guided framework. Extensive ablation studies verify the effectiveness of our proposed Pythonic-DSL and Scheduled Hill Climbing. We discuss the limitations and future directions of LLM-GS in Appendix J in detail, which includes the necessity of domain experts and LLMs, etc.
@@ Appendix A Extended Related work
Large language models and search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017], whose goal is to find programs that satisfy given examples, e.g., input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. The LLM can bootstrap program synthesis in the wake stage of library learning [Grand et al., 2024, Eberhardinger et al., 2023]. It is an open question whether the LLM can bootstrap the sleep stages. We utilize the description of the Karel tasks to bootstrap our scheduled HC. Additional information may be crucial for LLM assist bootstrapping and fine-tuning. This additional information may come from the execution [Chen et al., 2024], prompt design [Wang et al., 2023a, Wei et al., 2022, Wüst et al., 2024, Grand et al., 2024], and the task description in our work.
More on search methods and structured policy representations. There are several different search methods developed for programming by example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017] or DSLs with learned libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. Recent studies explore different representations for Reinforcement Learning policies toward interpretable and/or explainable. Together with various structured policy representations such as DSL programs [Andre and Russell, 2001, Verma et al., 2018, Silver et al., 2020], decision trees [Bastani et al., 2018, Kohler et al., 2024], state machines [Lin et al., 2024], and symbolic programs [Jiang and Luo, 2019, Qiu and Zhu, 2022, Delfosse et al., 2024]. Our method has the greatest bootstrapping ability for DSL programs but is difficult to extend to all kinds of representations. We designed our prompt by contrasting the policy representation (Karel DSL) with Python. As a result, our prompt in Section D.1 limits some of the functionalities of Python. We believe our framework is only suitable for tasks with task-solving procedures that could be described using languages. That said, it could be difficult to apply this framework to low-level control tasks, such as motor torque control [Qiu and Zhu, 2022].

### v3

@@ 2 Related work
Programmatic reinforcement learning (PRL). PRL represents RL policies using more structured and potentially more interpretable and generalizable representations, such as decision tree (Bastani et al., 2018), state machine (Inala et al., 2020; Koul et al., 2019; Lin et al., 2024), symbolic expression (Verma et al., 2018; 2019; Bhupatiraju et al., 2018; Landajuela et al., 2021), Logic programming language (Jiang and Luo, 2019), and program written in domain-specific language (DSL) (Andre and Russell, 2001; Silver et al., 2020; Sun et al., 2020; Zhu et al., 2019; Qiu and Zhu, 2022; Moraes and Lelis, 2024; Mariño et al., 2021; Trivedi et al., 2021; Liu et al., 2023; Carvalho et al., 2024). Trivedi et al. (2021); Liu et al. (2023) devised PRL tasks in the Karel domain (Pattis, 1981) and proposed learning embedding spaces of programs using variational autoencoders (Kingma and Welling, 2014) and then optimizing program embeddings using cross-entropy method or RL. Carvalho et al. (2024) achieve the state-of-the-art results in the Karel domain (Pattis, 1981) by searching in a programmatic space of AST structure. Despite the encouraging results and the generality of these methods, they are notoriously inefficient, requiring tens of millions of program executions to obtain task-solving programs. In this work, to devise an efficient PRL framework, we leverage the knowledge and the reasoning ability of LLMs to generate a set of initial programs to bootstrap search algorithms.
Large language models for code generation. With a remarkable ability to understand and generate natural languages and codes, LLMs have been widely adopted for code generation (Brown et al., 2020; Achiam et al., 2023; Nijkamp et al., 2022; Xu et al., 2022; Roziere et al., 2023). These works target general-purpose programming languages for software development, e.g. Python and C++, with abundant data on the internet; in contrast, we aim to solve PRL tasks that require writing programs in given domain-specific languages. Prior works exploring using LLMs to synthesize DSL programs via providing LLMs with DSL grammars and few-shot examples, hindsight relabeling, and prioritized experience replay (Wang et al., 2023a; Grand et al., 2024; Butt et al., 2024). While they focus on string/array transformations and abstract reasoning (Chollet, 2019), we leverage LLMs to synthesize DSL program policies to be executed in an RL environment and maximize the return.
@@ 3 Preliminary
Programmatic space. In the programmatic space, a program $P$ can be represented as an abstract syntax tree (AST), where each leaf node in the AST represents a program token (Carvalho et al., 2024). To obtain its neighborhood program, a node is sampled from its AST, and then replaced by a subtree randomly generated using the DSL’s production rules and sampling strategies.
Hill climbing (HC). This search algorithm climbs whenever there is a higher place, i.e., moves to a program having a higher episodic return. Given a program, HC generates its $k$ neighborhood programs from the search space. These programs are evaluated in the environment. If one of them has a higher episodic return than the initial program, it is set as the search center program, and then this process repeats. Otherwise, the algorithm halts and returns the best program evaluated. This search algorithm has a major weakness – it searches surrounding programs in narrow proximity; thus, the distance between initial programs and optimal programs could upper-bound its sample efficiency.
@@
4 Large language model guided search
@@ 4.2 Generating DSL programs with Pythonic-DSL strategy
Table 2: The Pythonic-DSL instruction. The Karel DSL is specified via a constrained version of Python with pre-defined functions, including Pythonic-DSL perceptions that allow the agent to observe, and Pythonic-DSL actions that enable the agent to interact. This Pythonic-DSL description is fed into the LLM via the system prompt so that it can generate Python code that can later be converted into Karel DSL. We will show that this approach outperforms the direct generation of DSL or Python programs in Table 3. The full prompt is presented in Section D.1.
@@ 4.3 Optimizing the program policy with scheduled hill climbing
where $n$ represents the number of evaluated programs, $k(n)$ denotes a function indicating the current number of neighborhood programs based on the number of evaluated programs $n$. This function’s logarithm is a linear combination of the logarithms of two hyperparameters, $K_{start}$ and $K_{end}$, which signify the initial and terminal numbers of neighbors to search, respectively. The variable $r(n)$ governs the linear ratio, which itself is a sinusoidal function that gradually increases from 00 to $1$ throughout the evaluation of a total of $N$ programs (see Appendix E). With this scheduler, the number of neighbors $k$ is a function of the execution budget used, growing from $K_{start}$ to $K_{end}$ on an exponential basis. We apply this scheduler to Hill Climbing and call this method Scheduled Hill Climbing (Scheduled HC).
@@ 5.1 Evaluation setup
Baselines. We compare our proposed framework with the following methods: LEAPS (Trivedi et al., 2021), HPRL (Liu et al., 2023), CEBS (Carvalho et al., 2024), and Hill Climbing (Carvalho et al., 2024) (the current state-of-the-art). Note that LEAPS and CEBS search programs in learned latent program space, as introduced in Appendix F.
Metrics. For a fair comparison in sample efficiency, we use the number of programs evaluated to represent the program-environment interaction budget, i.e. a method is more sample-efficient if it achieves a higher average episodic return on a fixed budget. Specifically, for each task, the number of task variances is $C$. Task variances arise from differences in the environment’s initial states. Each program evaluation means executing the program on all $C$ task variances to obtain an average return. We set the number of task variances $C=32$, the maximum number of program evaluation $N=10^{6}$, i.e., the interaction budget. Programs achieving an average return of $1.0$ are considered optimal.
@@ 5.2 LLM-guided search significantly improves the sample efficiency
Figure 4 presents the experimental result, indicating that our proposed LLM-GS surpasses all the existing methods by a large margin on almost all the tasks.
On the Karel set, our framework completely solves these tasks and exhibits the best sample efficiency. Also, LLM-GS performs extraordinarily well on Seeder, OneStroke, and DoorKey, which are among the hardest of the Karel-Hard set. We highlight our DoorKey result, with an efficiency improvement from not converging at even 1M (baselines) to converging within around 500K (ours). Existing PRL algorithms struggle at DoorKey due to the fact that it is a two-stage task with a sparse reward function. Therefore, most of the baselines cannot escape the local maxima and converge around $0.5$. Curiously, our method starts at an episodic return of $\approx 0.5$ from the beginning and converges quickly without being trapped by the local maxima. To analyze, we examine one output program from LLM shown in Figure 5. The LLM clearly understands that DoorKey is a two-stage task and thus generates a corresponding two-stage structured program, which is much easier for later search to find the optimal program. The full dialogue is presented in Appendix D.
We include a plot aggregating the performance across all ten tasks in Appendix H, a wall-clock time (real-time elapsed) evaluation in Appendix I, and a statistical significance test in Appendix J.
@@ 5.3 Ablation studies
To justify the effectiveness of our proposed Pythonic-DSL strategy, we conduct experiments on generating either Python or DSL programs with the same 8 seeds and present the results in Table 3. The acceptance rate calculates the ratio of executable DSL programs, while the best return refers to the highest episodic return among the legal programs. Our method achieves the highest acceptance rate on 9 / 10 tasks and the highest best return on 7 / 10 tasks. The complete prompts are presented in Appendix K.
@@
Figure 6: Comparing search algorithms. We compare the proposed Scheduled HC with latent space CEM, CEBS, HC-$250$, and Scheduled HC, and programmatic space HC-$k$$\in\{32,256,2048\}$ using LLM-initialized programs. Our method achieves comparatively good efficiency in DoorKey and performs best in CleanHouse, justifying the efficacy of our proposed search method.
@@
Figure 7: Comparing initializations. We compare initializing the search population using LLM-generated programs or randomly sampled programs, combined with our proposed SHC. The results demonstrate that starting from LLM-generated programs significantly improves efficiency and performance, especially in DoorKey, highlighting the effectiveness of our proposed framework.
@@
Scheduled HC variants. We ablate the components in the scheduled HC, i.e. logarithmic interpolation, sinusoidal schedule, and logarithmic ratio, and present the results in Section E.3.
@@ 5.4 Performance on novel tasks showcases the extensibility of LLM-GS
A key advantage of our proposed LLM-GS framework is its ability to adapt to new tasks using only text descriptions without requiring users to have any domain knowledge or familiarity with the DSL. To showcase this feature, we designed two novel Karel tasks, PathFollow and WallAvoider, as detailed in Section C.3. By simply replacing the task descriptions while keeping the Karel domain prompt, Python-to-DSL prompt, and hyperparameters the same, we compare our method with the best-performing baseline, HC. The results presented in Figure 9 show that by only changing task descriptions, LLM-GS still achieves improved sample efficiency and finds better program policies.
@@
Figure 8: Performance of PathFollow & WallAvoider. With only changing the task descriptions, LLM-GS surpasses the best-performing baseline HC on two novel tasks, highlighting the extensibility of LLM-GS.
(c) LavaGap
(d) PutNear
Figure 9: Performance of Minigrid LavaGap and PutNear. We compare LLM-GS and the best-performing baseline in Karel, i.e. HC, in the Minigrid domain. LLM-GS demonstrates better sample efficiency in both of the two tasks.
5.5 Performance on Minigrid domain verifies the adaptability of LLM-GS
To verify if our proposed LLM-GS framework can be adapted to another domain, we adopt three tasks in the Minigrid domain, LavaGap, PutNear, and RedBlueDoor. Unlike the case in Karel, the perception functions in our Minigrid DSL are parameterized, i.e., require an input parameter such as an object, as detailed in Appendix L. By adapting all aspects of the prompt from the Karel domain to the Minigrid version, including the domain prompt, the task descriptions, and the Python-to-DSL prompt, we evaluate our method against the strongest baseline, HC. The results presented in Figure 9 show that our proposed LLM-GS is more sample-efficient in the Minigrid domain.
@@ 5.6 Beyond initialization – a preliminary attempt on LLM revision
Given the encouraging results from LLM-initialized programs, one might be curious whether LLM itself without the help of any search algorithms can progressively revise and improve its programs once given feedback from the environment, e.g., reward. For completeness, we additionally conduct studies with the following LLM-based revision:
@@
Figure 10: LLM revision. The performance gain of LLM revision saturates within a few program revision rounds.
We conduct the LLM revision experiments on DoorKey for five revision rounds, each round with $32$ programs with $5$ seeds. The results are presented in Figure 10. We can see that the improvement quickly saturates within the first two rounds, and no significant gain is observed with more revision rounds. This indicates that only using LLM to revise the program may not be sufficient to solve the task. On the other hand, our proposed Scheduled HC can converge to an optimal program – and is “free,” i.e. no expensive API calls or “API usage has reached your notification threshold” emails. The details can be found in Appendix K.
@@ 6 Discussion
We propose the LLM-guided search (LLM-GS) that leverages the programming skills and common sense of LLM to bootstrap the efficiency of assumption-free, random-guessing search methods. We propose a Pythonic-DSL method to address LLMs’ inability to generate precise and grammatically correct programs. To improve the programs generated by LLMs, we design a search algorithm, Scheduled Hill Climbing, which can explore the programmatic search space to consistently improve the programs. The experimental results show the improved sample efficiency of LLM-GS. Extensive ablation studies verify the effectiveness of our proposed Pythonic-DSL method and Scheduled Hill Climbing. The limitations and future directions of LLM-GS are presented in Appendix M.
@@ Appendix A Extended Related work
Large language models and search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017], whose goal is to find programs that satisfy given examples, e.g., input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. Neural program synthesis methods [Lin et al., 2018, Sun et al., 2018, Desai et al., 2016, Raza et al., 2015, Wang et al., 2017, Zhong et al., 2023, Le et al., 2022, Parisotto et al., 2017, Balog et al., 2017] learn neural networks from data to synthesize programs intended by users. Grand et al. [2024], Eberhardinger et al. [2023] show that LLMs can bootstrap program synthesis in the wake stage of library learning; it is an open question whether the LLM can bootstrap the sleep stages. We utilize the description of the Karel tasks to bootstrap our scheduled HC. Additional information may be crucial for LLM assist bootstrapping and fine-tuning. This additional information may come from the execution [Chen et al., 2024], prompt design [Wang et al., 2023a, Wei et al., 2022, Wüst et al., 2024, Grand et al., 2024], and the task description in our work. More on search methods and structured policy representations. There are several different search methods developed for programming by example (PBE) [Gulwani, 2011, Feser et al., 2015, Polozov and Gulwani, 2015, Gulwani et al., 2017, Parisotto et al., 2017, Balog et al., 2017] or DSLs with learned libraries [Ellis et al., 2021, Grand et al., 2024, Eberhardinger et al., 2023, Wüst et al., 2024] for PBE, imitating policies [Bastani et al., 2018, Kohler et al., 2024], and object representations [Wüst et al., 2024]. Recent studies explore different representations for Reinforcement Learning policies toward interpretable and/or explainable. Together with various structured policy representations such as DSL programs [Andre and Russell, 2001, Verma et al., 2018, Silver et al., 2020], decision trees [Bastani et al., 2018, Kohler et al., 2024], state machines [Lin et al., 2024], and symbolic programs [Jiang and Luo, 2019, Qiu and Zhu, 2022, Delfosse et al., 2024]. Our method has the greatest bootstrapping ability for DSL programs but is difficult to extend to all kinds of representations. We designed our prompt by contrasting the policy representation (Karel DSL) with Python. As a result, our prompt in Section D.1 limits some of the functionalities of Python. We believe our framework is only suitable for tasks with task-solving procedures that could be described using languages. That said, it could be difficult to apply this framework to low-level control tasks, such as motor torque control [Qiu and Zhu, 2022].